OpenStudy (anonymous):

Which of the following numbers is not a potential rational zero of the polynomial function p(x) = x^4 − 9x^2 − 4x + 12 ?

OpenStudy (anonymous):

which numbers?

OpenStudy (anonymous):

(a) x = 1 (b) x = 2 (c) x = 3 (d) x = 4 (e) x = 5

OpenStudy (anonymous):

e) 5 Nothing multiplied by 5 can give you twelve or 1